# Variational Diffusion Posterior Sampling with Midpoint Guidance

### v1

@@ B.1 Derivation
In this section we derive the recursions verified by the first and second moments of the marginal distribution $\smash{\hat{\pi}^{\ell}_{0}}$ of the surrogate model (3.7) in the simplified setting of Example 3.2. We recall that in this specific example we assume that $q=\mathrm{N}(m,\Sigma)$ where $(m,\Sigma)\in\mathbb{R}^{d}\times\mathcal{S}^{++}_{d}$ and $g:x\mapsto\mathcal{N}(y;\mathrm{A}x,\sigma_{y}^{2}\mathrm{I}_{d_{y}})$.
@@
Since we are dealing with a Gaussian prior, the denoiser $\smash{m_{0|k}}$ can be computed in closed form for any $k\in\llbracket 1,n\rrbracket$. Using Bishop (2006, Eqn. 2.116), we have that
$q_{0|k}(x_{0}|x_{k})$ $\propto q(x_{0})q_{k|0}(x_{k}|x_{0})$
$=\mathcal{N}\big{(}x_{0};\Sigma_{0|k}\big{(}(\sqrt{\alpha_{k}}/v_{k})x_{k}+\Sigma^{-1}m\big{)},\Sigma_{0|k}\big{)},$
where $\Sigma_{0|k}\vcentcolon=((\alpha_{k}/v_{k})\mathrm{I}+\Sigma^{-1})^{-1}$. Hence,
$\smash{m_{0|k}}(x_{k})=\Sigma_{0|k}\big{(}(\sqrt{\alpha_{k}}/v_{k})x_{k}+\Sigma^{-1}m\big{)},$
and we assume in the remainder of this section that $\smash{m^{\theta}_{0|k}}=\smash{m_{0|k}}$. From the expression of $q_{0|k}(\cdot|x_{k})$ we can immediately derive the more general backward transitions $q_{\ell|k}(\cdot|x_{k})$ for $\ell\in\llbracket 1,k-1\rrbracket$ by noting that
$q_{\ell|k}(x_{\ell}|x_{k})=\int q_{\ell|0,k}(x_{\ell}|x_{0},x_{k})q_{0|k}(x_{0}|x_{k})\,\mathrm{d}x_{0}\,.$
From this, (A.3), (A.4), and the law of total expectation and covariance, it follows that $q_{\ell|k}(\cdot|x_{k})=\mathrm{N}(m_{\ell|k}(x_{k}),\Sigma_{\ell|k}(x_{k}))$, where
$m_{\ell|k}(x_{k})=m_{\ell|0,k}(\smash{m_{0|k}}(x_{k}),x_{k})\,,\quad\Sigma_{\ell|k}=\frac{\alpha_{\ell}(1-\alpha_{k}/\alpha_{\ell})^{2}}{(1-\alpha_{k})^{2}}\Sigma_{0|k}+v_{\ell|0,k}\,.$
@@
$p^{\theta}_{\ell|k}(\cdot|x_{k})=q_{\ell|0,k}(\cdot|\smash{m_{0|k}}(x_{k}),x_{k})=\mathrm{N}(m_{\ell|0,k}(\smash{m_{0|k}}(x_{k}),x_{k}),v_{\ell|0,k}\mathrm{I}_{d}),$ (B.1)
@@
For $k\in\llbracket 0,n\rrbracket$ we let $\smash{\hat{\pi}^{\ell}_{k}}$ denote the $x_{k}$ marginal of the surrogate model (3.7). We remind the reader that $\smash{\hat{\pi}^{\ell}_{n}}=\mathrm{N}(\mathrm{0}_{d},\mathrm{I}_{d})$. The marginals satisfy the recursion
$\smash{\hat{\pi}^{\ell}_{k}}(x_{k})=\int\smash{\hat{\pi}^{\ell}_{k|k+1}(x_{k}|x_{k+1})}\smash{\hat{\pi}^{\ell}_{k+1}}(x_{k+1})\,\mathrm{d}x_{k+1}\,,\quad k\in\llbracket 0,n-1\rrbracket\,.$
Since $\hat{g}^{\theta}_{k}(x_{k})=\mathcal{N}(y;\mathrm{A}\smash{m^{\theta}_{0|k}}(x_{k}),\sigma_{y}^{2}\mathrm{I}_{d_{y}})$ and $\smash{m^{\theta}_{0|k}}$ is linear in $x_{k}$, it is easily seen that $\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|x_{k+1})}$ is the density of a Gaussian distribution. Consequently, by definition (3.5) and the definition (A.5) of the bridge kernel, this is also the case for $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|x_{k+1})}$. Now assume that
$\smash{\hat{\pi}^{\ell}_{k+1}}(x_{k+1})$ $=\mathcal{N}(x_{k+1};\hat{\mu}^{\ell}_{k+1},\hat{\Sigma}^{\ell}_{k+1})\,,$ (B.2)
$\smash{\hat{\pi}^{\ell}_{k|k+1}(x_{k}|x_{k+1})}$ $=\mathcal{N}(x_{k};\mathrm{M}^{\ell}_{k|k+1}x_{k+1}+c^{\ell}_{k|k+1},\hat{\Sigma}^{\ell}_{k|k+1})\,,$ (B.3)
where $\mathrm{M}^{\ell}_{k|k+1}\in\mathbb{R}^{{d}\times{d}}$, $\hat{\Sigma}^{\ell}_{k|k+1}\in\mathcal{S}^{++}_{d}$, and $c^{\ell}_{k|k+1}\in\mathbb{R}^{{d}}$. Using the definition (A.5) of the bridge kernel we find that $\smash{\hat{\pi}^{\ell}_{k}}=\mathrm{N}(\hat{\mu}_{k},\hat{\Sigma}_{k})$, where
$\hat{\mu}^{\ell}_{k}$ $=\mathrm{M}^{\ell}_{k|k+1}\hat{\mu}^{\ell}_{k+1}+c^{\ell}_{k|k+1}\,,$
$\hat{\Sigma}^{\ell}_{k}$ $=\mathrm{M}^{\ell}_{k|k+1}\hat{\Sigma}^{\ell}_{k+1}\mathrm{M}^{\intercal}_{k|k+1}+\hat{\Sigma}^{\ell}_{k|k+1}\,.$
Iterating these updates until reaching $k=0$, starting from the initialization $\hat{\mu}^{\ell}_{n}=\mathrm{0}_{d}$ and $\hat{\Sigma}^{\ell}_{n}=\mathrm{I}_{d}$, yields the desired moments of the surrogate posterior $\smash{\pi}^{\ell}_{0}$. It now remains to show that the backward transition $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|x_{k+1})}$ writes in the form (B.3) and identify $\mathrm{M}^{\ell}_{k|k+1}$ and $c^{\ell}_{k|k+1}$.
First, we write the approximate potential in the form
$\hat{g}^{\theta}_{k}(x_{k})=\mathcal{N}(y;\hat{\mathrm{A}}_{k}x_{k}+b_{k},\sigma_{y}^{2}\mathrm{I}_{d_{y}}),$
@@
$\hat{\mathrm{A}}_{k}=(\sqrt{\alpha_{k}}/v_{k})\mathrm{A}\Sigma_{0|k},\quad b_{k}=\mathrm{A}\Sigma_{0|k}\Sigma^{-1}m\,.$
We also denote by $m^{\theta}_{\ell|k}(x_{k})$ the mean of the Gaussian distribution with density given by the DDPM transition $p^{\theta}_{\ell|k}(\cdot|x_{k})$ in (B.1). We have that
$m^{\theta}_{{\ell_{k}}|k+1}(x_{k+1})=\mathrm{H}_{{\ell_{k}}|k+1}x_{k+1}+h_{{\ell_{k}}|k+1}\,,$
@@
$\mathrm{H}_{{\ell_{k}}|k+1}$ :$\vcentcolon=\frac{\alpha_{{\ell_{k}}}(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})}{v_{\ell_{k}}v_{k+1}}\Sigma_{0|k}+\frac{\sqrt{\alpha_{k+1}}(1-\alpha_{{\ell_{k}}})}{v_{k+1}}\mathrm{I}_{d}\,,$ (B.4)
$h_{{\ell_{k}}|k+1}$ :$\vcentcolon=\frac{\sqrt{\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})}{v_{k+1}}\Sigma_{0|k}\Sigma^{-1}m\,.$ (B.5)
@@
$\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(x_{{\ell_{k}}}|x_{k+1})}=\mathcal{N}(x_{{\ell_{k}}};\widetilde{\mathrm{M}}^{\theta}_{{\ell_{k}}|k+1}x_{k+1}+\tilde{c}^{\theta}_{{\ell_{k}}|k+1},\Gamma_{{\ell_{k}}|k+1})\,,$
@@
$\Gamma_{{\ell_{k}}|k+1}$ :$\vcentcolon=\big{(}v^{-1}_{{\ell_{k}}|0,k+1}\mathrm{I}+\sigma_{y}^{-2}\hat{\mathrm{A}}_{{\ell_{k}}}^{\top}\hat{\mathrm{A}}_{{\ell_{k}}}\big{)}^{-1}\,,$
$\widetilde{\mathrm{M}}^{\theta}_{{\ell_{k}}|k+1}$ :$\vcentcolon=v^{-1}_{{\ell_{k}}|0,k+1}\Gamma_{{\ell_{k}}|k+1}\mathrm{H}_{{\ell_{k}}|k+1}\,,$
$\tilde{c}^{\theta}_{{\ell_{k}}|k+1}$ :$\vcentcolon=\Gamma_{{\ell_{k}}|k+1}\big{[}\sigma_{y}^{-2}\hat{\mathrm{A}}_{{\ell_{k}}}^{\top}(y-b_{{\ell_{k}}})+v^{-1}_{{\ell_{k}}|0,k+1}h_{{\ell_{k}}|k+1}\big{]}\,.$
@@
$\mathrm{M}^{\ell}_{k|k+1}$ $=\frac{\sqrt{\alpha_{k}/\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{k})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\widetilde{\mathrm{M}}^{\theta}_{{\ell_{k}}|k+1}+\frac{\sqrt{\alpha_{k+1}/\alpha_{k}}(1-\alpha_{k}/\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\mathrm{I}_{d}\,,$
$c^{\ell}_{k|k+1}$ $=\frac{\sqrt{\alpha_{k}/\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{k})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\tilde{c}^{\theta}_{{\ell_{k}}|k+1}\,.$

### v2

@@ B.1 Derivation
In this section we derive the recursions verified by the first and second moments of the marginal distribution $\smash{\hat{\pi}^{\ell}_{0}}$ of the surrogate model (3.6) in the simplified setting of Example 3.2. We recall that in this specific example we assume that $q=\mathrm{N}(\boldsymbol{m},\mathbf{\Sigma})$ where $(\boldsymbol{m},\mathbf{\Sigma})\in\mathbb{R}^{d}\times\mathcal{S}^{++}_{d}$ and $p({\boldsymbol{y}}|\cdot):\boldsymbol{x}\mapsto\mathcal{N}({\boldsymbol{y}};\mathbf{A}\boldsymbol{x},\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}})$.
@@
Since we are dealing with a Gaussian prior, the denoiser $\smash{\boldsymbol{m}_{0|k}}$ can be computed in closed form for any $k\in\llbracket 1,n\rrbracket$. Using Bishop (2006, Eqn. 2.116), we have that
$q_{0|k}(\boldsymbol{x}_{0}|\boldsymbol{x}_{k})$ $\propto q(\boldsymbol{x}_{0})q_{k|0}(\boldsymbol{x}_{k}|\boldsymbol{x}_{0})$
$=\mathcal{N}\big{(}\boldsymbol{x}_{0};\mathbf{\Sigma}_{0|k}\big{(}(\sqrt{\alpha_{k}}/v_{k})\boldsymbol{x}_{k}+\mathbf{\Sigma}^{-1}\boldsymbol{m}\big{)},\mathbf{\Sigma}_{0|k}\big{)},$
where $\mathbf{\Sigma}_{0|k}\vcentcolon=((\alpha_{k}/v_{k})\mathbf{I}+\mathbf{\Sigma}^{-1})^{-1}$. Hence,
$\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k})=\mathbf{\Sigma}_{0|k}\big{(}(\sqrt{\alpha_{k}}/v_{k})\boldsymbol{x}_{k}+\mathbf{\Sigma}^{-1}\boldsymbol{m}\big{)},$
and we assume in the remainder of this section that $\smash{\boldsymbol{m}^{\theta}_{0|k}}=\smash{\boldsymbol{m}_{0|k}}$. From the expression of $q_{0|k}(\cdot|\boldsymbol{x}_{k})$ we can immediately derive the more general backward transitions $q_{\ell|k}(\cdot|\boldsymbol{x}_{k})$ for $\ell\in\llbracket 1,k-1\rrbracket$ by noting that
$q_{\ell|k}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{k})=\int q_{\ell|0,k}(\boldsymbol{x}_{\ell}|\boldsymbol{x}_{0},\boldsymbol{x}_{k})q_{0|k}(\boldsymbol{x}_{0}|\boldsymbol{x}_{k})\,\mathrm{d}\boldsymbol{x}_{0}\,.$
From this, (A.3), (A.4), and the law of total expectation and covariance, it follows that $q_{\ell|k}(\cdot|\boldsymbol{x}_{k})=\mathrm{N}(\boldsymbol{m}_{\ell|k}(\boldsymbol{x}_{k}),\mathbf{\Sigma}_{\ell|k}(\boldsymbol{x}_{k}))$, where
$\boldsymbol{m}_{\ell|k}(\boldsymbol{x}_{k})=\boldsymbol{m}_{\ell|0,k}(\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k}),\boldsymbol{x}_{k})\,,\quad\mathbf{\Sigma}_{\ell|k}=\frac{\alpha_{\ell}(1-\alpha_{k}/\alpha_{\ell})^{2}}{(1-\alpha_{k})^{2}}\mathbf{\Sigma}_{0|k}+v_{\ell|0,k}\mathbf{I}_{{d}}\,.$
@@
$p^{\theta}_{\ell|k}(\cdot|\boldsymbol{x}_{k})=q_{\ell|0,k}(\cdot|\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k}),\boldsymbol{x}_{k})=\mathrm{N}(\boldsymbol{m}_{\ell|0,k}(\smash{\boldsymbol{m}_{0|k}}(\boldsymbol{x}_{k}),\boldsymbol{x}_{k}),v_{\ell|0,k}\mathbf{I}_{d}),$ (B.1)
@@
For $k\in\llbracket 0,n\rrbracket$ we let $\smash{\hat{\pi}^{\ell}_{k}}$ denote the $\boldsymbol{x}_{k}$ marginal of the surrogate model (3.6). We remind the reader that $\smash{\hat{\pi}^{\ell}_{n}}=\mathrm{N}(\mathbf{0}_{d},\mathbf{I}_{d})$. The marginals satisfy the recursion
$\smash{\hat{\pi}^{\ell}_{k}}(\boldsymbol{x}_{k})=\int\smash{\hat{\pi}^{\ell}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})}\smash{\hat{\pi}^{\ell}_{k+1}}(\boldsymbol{x}_{k+1})\,\mathrm{d}\boldsymbol{x}_{k+1}\,,\quad k\in\llbracket 0,n-1\rrbracket\,.$
Since $\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})=\mathcal{N}({\boldsymbol{y}};\mathbf{A}\smash{\boldsymbol{m}^{\theta}_{0|k}}(\boldsymbol{x}_{k}),\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}})$ and $\smash{\boldsymbol{m}^{\theta}_{0|k}}$ is linear in $\boldsymbol{x}_{k}$, it is easily seen that $\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\cdot|\boldsymbol{x}_{k+1})}$ is the density of a Gaussian distribution. Consequently, by definition (3.5) and the definition (A.1) of the bridge kernel, this is also the case for $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|\boldsymbol{x}_{k+1})}$. Now assume that
$\smash{\hat{\pi}^{\ell}_{k+1}}(\boldsymbol{x}_{k+1})$ $=\mathcal{N}(\boldsymbol{x}_{k+1};\hat{\boldsymbol{\mu}}^{\ell}_{k+1},\hat{\mathbf{\Sigma}}^{\ell}_{k+1})\,,$ (B.2)
$\smash{\hat{\pi}^{\ell}_{k|k+1}(\boldsymbol{x}_{k}|\boldsymbol{x}_{k+1})}$ $=\mathcal{N}(\boldsymbol{x}_{k};\mathbf{M}^{\ell}_{k|k+1}\boldsymbol{x}_{k+1}+\boldsymbol{c}^{\ell}_{k|k+1},\hat{\mathbf{\Sigma}}^{\ell}_{k|k+1})\,,$ (B.3)
where $\mathbf{M}^{\ell}_{k|k+1}\in\mathbb{R}^{{d}\times{d}}$, $\hat{\mathbf{\Sigma}}^{\ell}_{k|k+1}\in\mathcal{S}^{++}_{d}$, and $\boldsymbol{c}^{\ell}_{k|k+1}\in\mathbb{R}^{{d}}$. Using the definition (A.1) of the bridge kernel we find that $\smash{\hat{\pi}^{\ell}_{k}}=\mathrm{N}(\hat{\boldsymbol{\mu}}_{k},\hat{\mathbf{\Sigma}}_{k})$, where
$\hat{\boldsymbol{\mu}}^{\ell}_{k}$ $=\mathbf{M}^{\ell}_{k|k+1}\hat{\boldsymbol{\mu}}^{\ell}_{k+1}+\boldsymbol{c}^{\ell}_{k|k+1}\,,$
$\hat{\mathbf{\Sigma}}^{\ell}_{k}$ $=\mathbf{M}^{\ell}_{k|k+1}\hat{\mathbf{\Sigma}}^{\ell}_{k+1}\mathbf{M}^{\intercal}_{k|k+1}+\hat{\mathbf{\Sigma}}^{\ell}_{k|k+1}\,.$
Iterating these updates until reaching $k=0$, starting from the initialization $\hat{\boldsymbol{\mu}}^{\ell}_{n}=\mathbf{0}_{d}$ and $\hat{\mathbf{\Sigma}}^{\ell}_{n}=\mathbf{I}_{d}$, yields the desired moments of the surrogate posterior $\smash{\pi}^{\ell}_{0}$. It now remains to show that the backward transition $\smash{\hat{\pi}^{\ell}_{k|k+1}(\cdot|\boldsymbol{x}_{k+1})}$ writes in the form (B.3) and identify $\mathbf{M}^{\ell}_{k|k+1}$ and $\boldsymbol{c}^{\ell}_{k|k+1}$.
First, we write the approximate likelihood in the form
$\hat{p}^{\theta}_{k}({\boldsymbol{y}}|\boldsymbol{x}_{k})=\mathcal{N}({\boldsymbol{y}};\hat{\mathbf{A}}_{k}\boldsymbol{x}_{k}+\boldsymbol{b}_{k},\sigma_{\boldsymbol{y}}^{2}\mathbf{I}_{d_{\boldsymbol{y}}}),$
@@
$\hat{\mathbf{A}}_{k}=(\sqrt{\alpha_{k}}/v_{k})\mathbf{A}\mathbf{\Sigma}_{0|k},\quad\boldsymbol{b}_{k}=\mathbf{A}\mathbf{\Sigma}_{0|k}\mathbf{\Sigma}^{-1}\boldsymbol{m}\,.$
We also denote by $\boldsymbol{m}^{\theta}_{\ell|k}(\boldsymbol{x}_{k})$ the mean of the Gaussian distribution with density given by the DDPM transition $p^{\theta}_{\ell|k}(\cdot|\boldsymbol{x}_{k})$ in (B.1). We have that
$\boldsymbol{m}^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{k+1})=\mathbf{H}_{{\ell_{k}}|k+1}\boldsymbol{x}_{k+1}+\boldsymbol{h}_{{\ell_{k}}|k+1}\,,$
@@
$\mathbf{H}_{{\ell_{k}}|k+1}$ :$\vcentcolon=\frac{\alpha_{{\ell_{k}}}(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})}{v_{\ell_{k}}v_{k+1}}\mathbf{\Sigma}_{0|k}+\frac{\sqrt{\alpha_{k+1}}(1-\alpha_{{\ell_{k}}})}{v_{k+1}}\mathbf{I}_{d}\,,$ (B.4)
$\boldsymbol{h}_{{\ell_{k}}|k+1}$ :$\vcentcolon=\frac{\sqrt{\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{{\ell_{k}}})}{v_{k+1}}\mathbf{\Sigma}_{0|k}\mathbf{\Sigma}^{-1}\boldsymbol{m}\,.$ (B.5)
@@
$\smash{\hat{\pi}^{\theta}_{{\ell_{k}}|k+1}(\boldsymbol{x}_{{\ell_{k}}}|\boldsymbol{x}_{k+1})}=\mathcal{N}(\boldsymbol{x}_{{\ell_{k}}};\widetilde{\mathbf{M}}^{\theta}_{{\ell_{k}}|k+1}\boldsymbol{x}_{k+1}+\tilde{\boldsymbol{c}}^{\theta}_{{\ell_{k}}|k+1},\mathbf{\Gamma}_{{\ell_{k}}|k+1})\,,$
@@
$\mathbf{\Gamma}_{{\ell_{k}}|k+1}$ :$\vcentcolon=\big{(}v^{-1}_{{\ell_{k}}|0,k+1}\mathbf{I}+\sigma_{{\boldsymbol{y}}}^{-2}\hat{\mathbf{A}}_{{\ell_{k}}}^{\top}\hat{\mathbf{A}}_{{\ell_{k}}}\big{)}^{-1}\,,$
$\widetilde{\mathbf{M}}^{\theta}_{{\ell_{k}}|k+1}$ :$\vcentcolon=v^{-1}_{{\ell_{k}}|0,k+1}\mathbf{\Gamma}_{{\ell_{k}}|k+1}\mathbf{H}_{{\ell_{k}}|k+1}\,,$
$\tilde{\boldsymbol{c}}^{\theta}_{{\ell_{k}}|k+1}$ :$\vcentcolon=\mathbf{\Gamma}_{{\ell_{k}}|k+1}\big{[}\sigma_{{\boldsymbol{y}}}^{-2}\hat{\mathbf{A}}_{{\ell_{k}}}^{\top}({\boldsymbol{y}}-\boldsymbol{b}_{{\ell_{k}}})+v^{-1}_{{\ell_{k}}|0,k+1}\boldsymbol{h}_{{\ell_{k}}|k+1}\big{]}\,.$
@@
$\mathbf{M}^{\ell}_{k|k+1}$ $=\frac{\sqrt{\alpha_{k}/\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{k})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\widetilde{\mathbf{M}}^{\theta}_{{\ell_{k}}|k+1}+\frac{\sqrt{\alpha_{k+1}/\alpha_{k}}(1-\alpha_{k}/\alpha_{{\ell_{k}}})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\mathbf{I}_{d}\,,$
$\boldsymbol{c}^{\ell}_{k|k+1}$ $=\frac{\sqrt{\alpha_{k}/\alpha_{{\ell_{k}}}}(1-\alpha_{k+1}/\alpha_{k})}{1-\alpha_{k+1}/\alpha_{{\ell_{k}}}}\tilde{\boldsymbol{c}}^{\theta}_{{\ell_{k}}|k+1}\,.$